OpenStudy (anonymous):

Jim Abbott purchased a $60,000 RV with a 40% markup on selling price what was the amount of the dealers markup what was the dealers original cost

jimthompson5910 (jim_thompson5910):

x = original cost (in dollars) x + 40% of x = final price x + 0.40x = 60,000 1x + 0.40x = 60,000 1.40x = 60,000 x = ???

jimthompson5910 (jim_thompson5910):

I'll let you finish up